OpenStudy (anonymous):

Determine the % composition of ammonium nitrate: NH4NO3

OpenStudy (anonymous):

precent composition i think. What i typed is the question(this time no typos)

OpenStudy (chmvijay):

can you tell molecular weigh of NH4NO3

OpenStudy (anonymous):

80.052

OpenStudy (chmvijay):

how many diffrernt elements are present in NH4NO3 can you tell

OpenStudy (anonymous):

4

OpenStudy (chmvijay):

what are those ??

OpenStudy (anonymous):

N - nitrogen H4 - hydrogen N - nitrogen O3 - oxygen

OpenStudy (chmvijay):

can you tell the atomic mass of N, O and H

OpenStudy (anonymous):

14.007 - N 15.999 - O 1.008 - H

OpenStudy (chmvijay):

now in 80 gram of NH4NO3 you have 2*14 gramof N right then for 100 gram =2*14*100 /80=

OpenStudy (chmvijay):

can you calculate this similarly do for the O and H too

OpenStudy (anonymous):

2 x 14 x 100 / 80 = 35 2 x 16 x 100 / 80 = 40 2 x 1 x 100 / 80 = 2.5

OpenStudy (chmvijay):

you should not take 2 for all that is 3 for O and 4 for hydrogen :)

OpenStudy (anonymous):

oh so 2 x 14 x 100 / 80 = 35 3 x 16 x 100 / 80 = 60 4 x 1 x 100 / 80 = 5

OpenStudy (chmvijay):

yaaa now total is 100 % right the compostion of NH4 NO# is 35 % N ,60% O and 5 % H